Today during the recovery did several changes to complete the OMC AA
- Small change in B5 QD2 offset in vertical (by 50um if I remember well)
- Add offset for the B1s QD2 Q quadrature that is used in loop
- Increase gain of B1s QD2 Q by factor 20, to match what was used for the I quadrature in the past
- Close dither controls on OMC in LN2, and increase gains (especially in TX to match the 20 times larger RF control signal)
After this while checking that the OMC AA dither doesn't misbehave (and lowering its gain by a factor 2), did a couple of short tests.
11:38:00 (2min) PRCL offset adjusted to bring frequency noise coupling to B2 to zero.
Made also by mistake small SRCL offset adjusments.
Figure 1 shows that the PRCL offset adjustment, also slightly incraesed the 12MHz on B4.
Figure 2 shows that the test of PRCL offsets 10 days ago also had a small impact on the 12MHz magnitude on B4.
11:52 (2 min) reference time
11:55 (2 min) B2 8MHz FF with 1e-2 weight
11:58 (2 min) B2 8MHz FF with 3e-2 weight
12:01 (2 min) B2 8MHz FF with -1e-2 weight
12:05 (2 min) B2 8MHz FF with -3e-2 weight
12:08 (2 min) reference time
12:10:30 (2 min) B2 8MHz FF with -0.5e-2 weight